### v1

# Merging of Kolmogorov-Arnold networks trained on disjoint datasets

Abstract
Training on disjoint datasets can serve two primary goals: accelerating data processing and enabling federated learning. It has already been established that Kolmogorov-Arnold networks (KANs) are particularly well suited for federated learning and can be merged through simple parameter averaging. While the federated learning literature has mostly focused on achieving training convergence across distributed nodes, the present paper specifically targets acceleration of the training, which depends critically on the choice of an optimisation method and the type of the basis functions. To the best knowledge of the authors, the fastest currently-available combination is the Newton-Kaczmarz method and the piecewise-linear basis functions. Here, it is shown that training on disjoint datasets (or disjoint subsets of the training dataset) can further improve the performance. Experimental comparisons are provided, and all corresponding codes are publicly available.
Keywords: Newton-Kaczmarz method, piecewise-linear functions, training on disjoint datasets.
1  Introduction
Since the publication of Liu et al. [1], which advocated spline-based univariate functions together with the ‘Adam’ or LBFGS optimisers for the training, these design choices have become dominant in the recent literature on Kolmogorov-Arnold networks (KANs). Several works have proposed alternative basis functions, aimed at reducing the computational overhead relative to neural networks, such as radial or alternative polynomial bases. However, these approaches remain conceptually similar — they rely on smooth basis functions [^1] Term ‘smooth’ is used loosely, implying the existence of some number of continuous derivatives; formally, the term of course means belonging to $C^{\infty}$ class. and typically require a relatively large number of calculations per each underlying function of KAN. Earlier alternatives, predating [1], include ReLU-type and piecewise-linear basis functions [2, 3], which were observed to be faster, but most subsequent research followed [1] as the principal reference point. It is important to emphasise that KANs were not invented in 2024 and originate from the pioneering work of Igelnik and Parikh [4] published in 2003. A historical overview of KANs’ development can be found in e.g. [5] and references therein.
A number of recent studies have evaluated the federated learning performance of KANs relative to multilayer perceptrons (MLPs) [6, 7, 8, 9]. These works also adopt the spline-based architectures and various optimisation methods. While improved accuracy for KANs is consistently reported, the computational cost is significantly higher. For example, the authors of [6] report training times of $314\,\text{s}$ for their KAN variant vs. $2.35\,\text{s}$ for an MLP on the same task. In the present paper, it is shown that such difference is not intrinsic to the KAN model class; rather, it arises from particular choices of basis functions and optimisation algorithms. For the piecewise-linear basis functions and the Newton-Kaczmarz optimisation scheme, this disparity is reversed, and highly-efficient KANs’ training is achieved while maintaining comparable accuracy. Moreover, in the present paper, for the first time, it is shown that the application of the federated-learning idea, i.e. training on disjoint subsets of a training dataset and merging the models, can significantly accelerate the training and represents a form of parallelisation of the training.
Although basis functions with continuous derivatives are essential for some applications, such as physics-informed modelling, they are not universally required. Moreover, the basis used during the training does not need to coincide with the basis used in the final model. In particular, a model may be trained efficiently using a piecewise-linear representation and subsequently upgraded to splines or other smooth bases with only minor post-training, thereby preserving accuracy while substantially reducing the training time. Thus, selecting piecewise-linear functions does not represent a move towards a more primitive model; instead, it enables faster optimisation while retaining full descriptive capabilities. The latter is further elaborated in [5].
2  Approach
A Kaczmarz-based training method for KANs with piecewise-linear basis functions was introduced by the authors in 2020 in a preprint later published as [3]; the approach was generalised to the Newton-Kaczmarz training method for KANs with arbitrary basis functions in 2023 in a preprint later published as [5]. The reader is referred to the above publications for details; however, for completeness of the paper, a brief summary is provided below, assuming that the reader is familiar with KANs.
The basic building block of KAN is an operator that converts $n$-dimensional vector $\boldsymbol{y}$ into scalar $\hat{z}$. It can be called the discrete Urysohn operator or a generalised additive model:
$\hat{z}=\sum_{j=1}^{n}g_{j}\left(y_{j}\right),$ (1)
where $g_{j}$ are some functions. A layer of KAN consists of a set of such blocks, forming $m$-dimensional vector $\hat{\boldsymbol{z}}$, such that
$\hat{z}_{i}=\sum_{j=1}^{n}g_{i,j}\left(y_{j}\right),\quad i=1,2,\ldots,m.$ (2)
Thus, a layer converts vector $\boldsymbol{y}$ into vector $\hat{\boldsymbol{z}}$. KAN is a chain of such layers, where the outputs of the previous layer are the inputs for the next layer.
Here, functions $g_{i,j}$ are taken to be piecewise-linear. Each argument $y_{j}$ lies within a single linear segment. For uniformly-spaced nodes with inter-nodal distance $\Delta y$, the interpolation is trivially computed via
$s=1+\left(y_{j}-y_{\mathrm{min}}\right)/\Delta y,\qquad k=\lfloor s\rfloor,\qquad f=s-k,$ (3)
$g_{i,j}\left(y_{j}\right)=\left(1-f\right)G_{i,j,k}+fG_{i,j,k+1},$ (4)
where $k$ is the index of the left node, $f$ is the relative offset from that node, and $G_{i,j,k}$ are the model’s parameters (values of $g_{i,j}$ at the nodes).
Training of KAN consists in updating parameters layer-by-layer for each record of the training dataset. To update one layer, a ‘desired’ (i.e. ‘target’) vector $\boldsymbol{z}$ is required. One can execute such update by performing one ‘damped’ step of a gradient descend, applied to layer loss function $L=\left\lVert\boldsymbol{z}-\boldsymbol{\hat{z}}\right\rVert^{2}$, resulting in
$G_{i,j,k}^{\mathrm{new}}=G_{i,j,k}^{\mathrm{old}}+\mu\left(z_{i}-\hat{z}_{i}\right)\left(1-f\right),$ (5)
$G_{i,j,k+1}^{\mathrm{new}}=G_{i,j,k+1}^{\mathrm{old}}+\mu\left(z_{i}-\hat{z}_{i}\right)f,$ (6)
where $k$ and $f$ depend on index $j$ according to equation (3), multiplier $\mu$ is the scaling factor, which depends on all layers in the original form of the method.
For the outer layer of KAN, the target vector is given in the training dataset. For some given layer, if target vector $\boldsymbol{z}$ is known, then one can calculate target vector $\boldsymbol{y}$ for the preceding layer simply by the chain rule for the derivatives:
$\Delta\boldsymbol{y}=J^{\mathrm{T}}\Delta\boldsymbol{z},$ (7)
where $\Delta\boldsymbol{z}=\boldsymbol{z}-\boldsymbol{\hat{z}}$ and $\Delta\boldsymbol{y}=\boldsymbol{y}-\boldsymbol{\hat{y}}$ are the residuals; matrix $J$ consists of partial derivatives, such that $J_{i,j}=\partial g_{i,j}/\partial y_{j}$, which are trivial to compute for piecewise-linear functions.
3  Merging of the models
The described approach iterates record-by-record through the training dataset. The hypothesis of the present paper is that such training can be parallelised using the algorithm as follows.
A number of equal-size subsets (batches) are formed using the records from the training dataset. Each batch should be sufficiently representative of the modelled system, i.e. it is implied that the algorithm will be used for large data. Hence, one should choose the number of batches and the batch size.
For each batch, an exact copy of the model (KAN) is created and trained on the corresponding batch (one pass through the data). The models are trained in parallel. After the training is finished, the models are merged into a single model simply by computing the mean value of every parameter.
The above steps (forming batches, creating copies of the model, training copies, merging copies) are iteratively repeated until the selected convergence criteria is reached. The performance of this approach should of course depend on the number of batches and on the batch size. In contrast to federated learning, these parameters may be chosen arbitrarily.
Using this approach, processing of the records from the training dataset should be parallelisable almost ideally, with minor losses for model coping/merging. However, there will be some loss of accuracy due to the merging, and one may expect increasing error with the number of batches. Therefore, there should be an optimal number of batches. Thus, the computational experiments serve two completely separate purposes: the first is to establish the scaling of the implementation of the algorithm (i.e. how much time is lost for coping/merging), the second is to find how the accuracy drops with the number of batches.
The loss of accuracy due to the merging is related to the parallel models all being equal in the beginning of the training on the batches. However, in the training algorithm, the model’s update depends on the current model’s parameters. Thus, in the sequential case, each update is performed for the up-to-date model, while in the parallel case, threads starting from the second work with the ‘old’ model.
4  Experimental part performed on laptop — strong scaling
4.1  Baseline
The single-threaded implementation of the approach described in section 2 has already been extensively tested in the previous publication by the authors. It has been shown that it significantly outperforms state-of-the-art MATLAB’s neural networks. However, for completeness of the present paper, it is worth re-establishing a baseline result — showing a brief comparison between the considered single-threaded (sequential) implementation and MATLAB’s neural networks.
Tests are conducted on a CPU (Intel(R) Core(TM) Ultra 7 155H, 16 cores) and on a GPU (Nvidia GeForce RTX 4050 Laptop). Obviously, one cannot objectively compare a CPU code with a GPU code, as they run on completely different hardware. However, the authors believe that their setup is representative of a typical modern researcher’s laptop, and it is beneficial to see the CPU/GPU difference in this particular case, i.e. whether performance gain of MATLAB’s neural networks when switching to a GPU is sufficient to overtake the authors KANs’ implementation on such particular hardware.
For the tests, synthetic data is generated. The inputs are elements of $4\times 4$ matrices and the outputs are the determinants of the matrices. The training dataset contains $100\mathrm{K}$ records; the validation dataset contains $20\mathrm{K}$ records. Random data is generated each execution of the code. The accuracy metric is the Pearson correlation coefficient (in percent) between the real and the modelled output.
This is just a demonstrative example. Therefore, only one configuration of MATLAB’s neural network is used, with $5$ hidden layers, each consisting of $10$ neurons (i.e. the total of $621$ parameters per model); only one KAN configuration in the authors’ code is used, with $70$ addends (classical two-layer architecture), $3$ points per inner function, and $25$ points per outer function (i.e. the total of $5,110$ parameters per model). Readers who are interested in a more detailed comparison with varying MLP and KAN architectures are referred to the previous paper [5].
The results are shown in table 1, where it is seen that MLPs are outperformed by the authors’ implementation of KANs. When switching to a GPU, MATLAB’s neural networks are approximately $3$ times faster (on the particular tested hardware), but still have slightly lower accuracy and longer training time than KANs.
MLP, CPU MLP, GPU KAN
Pearson (%) $94.6\pm 1.4$ $95.0\pm 0.9$ $97.1\pm 0.1$
time (s) $32.7\pm 0.9$ $8.6\pm 0.3$ $5.9\pm 0.6$
Table 1: Baseline accuracy and performance tests for modelling of the determinants of random $4\times 4$ matrices conducted for MATLAB’s neural networks and the authors’ KANs’ implementation (single-threaded, C++ implementation for CPU). The values are expectations and standard deviations for 88 executions.
4.2  Scaling test
In this section, the main hypothesis of the paper that parallelisation can be implemented by training on disjoint batches across multiple threads is benchmarked. The strong scaling is tested: the total computational work is fixed, which is the total number of processed records of the training dataset, while the work is equally split between the threads. The batch size and the number of threads are varied. As usual, the speedup is measured as $S=t(1)/t(n)$, where $t(1)$ is the time required to complete the work on $1$ thread and $t(n)$ is the time required to complete the same work on $n$ threads.
In the single-threaded setting for predicting determinants of $4\times 4$ matrices, approximately $10$ passes through the training dataset (epochs) are required to reach the Pearson correlation of at least $97\%$. This corresponds to processing $10^{6}$ records, and this value is used as the fixed workload for all parallel configurations. The number of threads is varied from $1$ to $6$ (the particular Intel CPU has $6$ performance ‘fast’ cores and $10$ efficiency ‘slow’ cores). Batch sizes range from $50\mathrm{K}$ down to $3\mathrm{K}$ records. For each combination of thread count and batch size, the required number of model merges (rounds) is chosen such that the total number of processed records remains exactly $1\mathrm{M}$.
The results are shown in table 2, where it is seen that the scaling is sufficiently close to linear, given that the experiments are performed on a Windows laptop (i.e. there is little control over background activity, and the hardware is not as optimised for parallel execution as the HPC hardware). The speedup on $4$ threads is between $3.7$ and $4.1$, depending on the batch size, while the speedup on $6$ threads is between $4.5$ and $4.9$.
Under ideal scaling, the accuracy would remain unchanged; however, the merging of the models leads to a computational error, and the actual value drops from $96.8\%$ for $1$ thread to $94.5\%$ for $6$ threads. Such loss in accuracy can be compensated of course by increasing the number of rounds. For example, using $6$ threads, $20$ rounds, and the batch size of $16,666$ (twice larger computational work than in the table) yields the Pearson correlation of $97.0\%$, while completing in $2.22$ seconds — $2.4$ times faster than for $1$ thread.
threads rounds batch size Pearson (%) time (s)
1 10 100,000 96.8 5.4
2 10 50,000 96.7 2.58
2 20 25,000 96.6 2.44
2 40 12,500 96.3 2.63
3 10 33,333 96.6 1.83
3 20 16,666 96.0 1.86
3 40 16,666 96.1 1.79
4 10 25,000 96.1 1.32
4 20 12,500 95.7 1.42
4 40 6,250 95.8 1.46
5 10 20,000 95.5 1.21
5 20 10,000 94.9 1.27
5 40 5,000 95.1 1.20
6 10 16,666 94.5 1.10
6 20 8,333 94.4 1.12
6 40 4,166 94.6 1.19
Table 2: Accuracies and execution times of the code on a laptop for the example of matrix determinants — strong scaling test.
4.3  Second example
The second example is not strictly a scaling test, as the computational workload varies, but it serves as a benchmark of parallel training to a fixed accuracy. The training is done on synthetic data, where the outputs are the areas of triangular faces of randomly-generated tetrahedra ($4$ faces), and the inputs are the coordinates of the vertices of the tetrahedra in 3D ($4$ vertices, $3$ coordinates for each). Number of the training records is $500\mathrm{K}$, number of the validation records is $50\mathrm{K}$. Three-layered KANs are trained, the first layer consisting of $70$ operators with $3$ points per function, the second layer consisting of $30$ operators with $18$ points per function, the third layer consisting of $4$ operators with $22$ points per function (i.e. the total of $42,960$ model parameters). MATLAB’s neural networks are used for the comparison with $4$ hidden layers and $80$ neurons per layer ($20,804$ model parameters). The results are shown in table 3, where it is seen that when trained to the same accuracy, parallel implementation of KANs is about $2.8$ times faster than the sequential and two orders of magnitude faster than MLP training on CPU.
Pearson (%) for outputs time (s)
MLP, CPU $98.6\pm 0.7$ $98.2\pm 0.9$ $98.9\pm 0.3$ $97.9\pm 0.4$ $1126.4\pm 69.5$
MLP, GPU $98.7\pm 0.5$ $98.3\pm 0.7$ $98.2\pm 0.8$ $98.3\pm 0.9$ $76.9\pm 1.2$
KAN sequential $8.6\pm 0.6 99$$9.0\pm 0.4$ $98.1\pm 0.2$ $98.5\pm 0.3$ $29.2\pm 3.2$
KAN parallel, 6 threads $98.5\pm 0.9$ $98.1\pm 0.7$ $98.6\pm 0.3$ $98.3\pm 0.4$ $10.5\pm 1.3$
Table 3: Baseline accuracy and performance tests for modelling of the areas of faces of random tetrahedrons conducted for MATLAB’s neural networks and the authors’ KANs’ implementation.
5  Experimental part performed on cluster — weak scaling
To assess the scalability of the code, the tests were run on an HPC cluster. All jobs were run on a single node equipped with two Intel(R) Xeon(R) Platinum 8358 CPUs ($32$ cores per CPU) and $256\mathrm{GB}$ RAM. The tests consisted in training KANs on datasets with the inputs being elements of $4\times 4$ or $5\times 5$ matrices and the outputs being the determinants of the matrices.
In this section, the weak scaling is tested: the number of threads was kept proportional to the total number of passes through the data (epochs), i.e. the total computational work. Obviously, in the case of ideal (linear) scaling, the execution times should be the same. As usual, the efficiency is measured as $E=t(1)/t(n)$, where $t(1)$ is the time required to complete work on $1$ thread and $t(n)$ is the time require to complete work that is $n$ times larger on $n$ threads.
The results are shown in table 4, where it is seen that the scaling is very close to linear. For the example of $4\times 4$ matrix determinants, the efficiency is above $93\%$ when up to $32$ threads are used. For the example of $5\times 5$ matrix determinants, the efficiency is above $95\%$ when up to $16$ threads are used and only drops down to $83\%$ when $32$ threads are used (and even further for $64$ threads). Such noticeable decline in the latter case is due to a large number of model synchronisations and large model sizes for $5\times 5$ example.
In the example of $4\times 4$ matrix determinants, the datasets consisted of $100\mathrm{K}$ training records, the models consisted of $70$ addends, with $3$ points per inner function and $25$ points per outer function (i.e. the total of $5,110$ parameters per model). The batch size was taken to be $100\mathrm{K}$ records, i.e. the entire training dataset. The jobs consisted of $10$ batches, i.e. $10$ synchronisations (of $n$ models, where $n$ is the number of threads) are made within each execution of the code.
In the example of $5\times 5$ matrix determinants, the datasets consisted of $10\mathrm{M}$ training records, the models consisted of $160$ addends, with $3$ points per inner functions and $30$ points per outer function (i.e. the total of $16,800$ parameters per model). The batch size was taken to be $50\mathrm{K}$ records. The jobs consisted of $200$ batches.
threads 11 22 44 88 $16$ $32$ $64$
time (s), $4\times 4$ $6.64$ $6.80$ $6.82$ $6.82$ $7.11$ $7.04$ $7.79$
time (s), $5\times 5$ $217.02$ $220.24$ $226.07$ $222.82$ $227.68$ $260.99$ $282.96$
Table 4: Execution times of the code on HPC cluster for examples of matrix determinants — weak scaling test. The values are averages of three programme runs.
6  Concurrent-serial alternations of Newton-Kaczmarz
The conceptual description of the Newton-Kaczmarz algorithm provided in section 2 highlights its alternating concurrent-sequential structure. Each layer maps an input vector to an output vector through multiple independent operations that may be executed concurrently. However, the outputs of all concurrent computations must be synchronised before proceeding to the next layer. The backward updates follow the same pattern — a chain of potentially concurrent steps separated by unavoidable sequential dependencies.
These serial steps between independent computations represent a significant obstacle for concurrent training. Neither CPU nor GPU kernels are well suited for such patterns. This limitation can be illustrated using the classical vector-input-scalar-output two-layer KAN:
$\hat{z}=\sum_{j=1}^{n}A_{j},\qquad A_{j}=g_{j}\left(\sum_{l=1}^{p}f_{j,l}\left(x_{l}\right)\right),$ (8)
where $A_{j}$ stands for one addend, and each addend is itself a small tree mapping inputs $x_{l}$ to a scalar. All $n$ addends must be computed in order to obtain the residual $\Delta z$, and these computations can be executed concurrently. This residual, divided evenly among the addends and scaled by the learning rate, is the only quantity required for updating all functions $g_{j}$ and $f_{j,l}$. In the piecewise linear model, evaluating a function requires only two operations (one division and one multiplication), which is far too little computation for concurrency on CPUs or GPUs to provide any advantage. The synchronisation required after all $n$ addends are computed often takes longer than the time saved through parallelism.
However, this pattern is well suited for customised hardware, such as field-programmable gate arrays (FPGAs) or application-specific integrated circuits (ASICs). A CPU-based simulation demonstrating a custom synchronisation barrier designed by the authors is available online [^2] https://github.com/andrewpolar/detfast, but it serves only as a conceptual demonstration; synchronising at each training record introduces substantial overhead. The example implemented in the code is training a model for determinates of $5\times 5$ matrices. Training MLPs for the latter requires several hours [5], whereas the CPU-based KAN code completes it in a few minutes. An FPGA or ASIC implementation could reduce this further to seconds, because each addend can be mapped to a dedicated hardware processing block. In such customised hardware, concurrency is not emulated through threads or kernels but is realised through physical parallelism, where synchronization and data flow are governed directly by wiring rather than by software-controlled barriers.
6.1  Pretraining
Structure (8) motivates a pretraining strategy, in which groups of addends are trained concurrently over the entire dataset to predict an equal share of target $z$. For the example of determinants of $5\times 5$ matrices and a dataset of $10\mathrm{M}$ records, such pretraining produces a model with the Pearson correlation near $70\%$, providing a strong initialisation for the main training phase. When pretraining is combined with the concurrent-serial alternations, the full training process for $5\times 5$ determinants completes in just over $2$ minutes on the CPU used for the baseline examples.
7  Conclusions
The results demonstrate that KANs’ training can indeed be accelerated by iteratively training copies of the model on disjoint datasets (batches) and merging them. The speedup ratios between sequential and concurrent KAN training typically fall in the range of 3344, meaning that acceleration is possible, but there are some limits.
At the same time, even the sequential KAN implementation by the authors consistently outperforms MLPs on CPUs and, remarkably, also on a GPU. While direct CPU-GPU comparisons are formally inappropriate due to differences in hardware, these results are included for reference, taking into account that the GPU used for the benchmarking can be considered to be a typical laptop GPU (it has been noticed by the authors that some researchers have unrealistic expectations from GPU implementations).
Since methodology adopted in the authors’ code is FPGA-friendly, the authors believe that this direction is a perspective research capable to reduce hours into seconds of processing time for very large models and datasets.
MATLAB scripts [^3] http://openkan.org/downloadskan/matlabscripts.zip, sequential implementation [^4] https://github.com/andrewpolar/kankan-8, and parallel implementation [^5] https://github.com/andrewpolar/meanmodel are available online.

### v2

@@
-Merging of Kolmogorov-Arnold networks trained on disjoint datasets
+Enhancing Newton-Kaczmarz training of Kolmogorov-Arnold networks through concurrency
 Abstract
-Training on disjoint datasets can serve two primary goals: accelerating data processing and enabling federated learning.
-It has already been established that Kolmogorov-Arnold networks (KANs) are particularly well suited for federated learning and can be merged through simple parameter averaging.
-While the federated learning literature has mostly focused on achieving training convergence across distributed nodes, the present paper specifically targets acceleration of the training, which depends critically on the choice of an optimisation method and the type of the basis functions.
-To the best knowledge of the authors, the fastest currently-available combination is the Newton-Kaczmarz method and the piecewise-linear basis functions.
-Here, it is shown that training on disjoint datasets (or disjoint subsets of the training dataset) can further improve the performance.
-Experimental comparisons are provided, and all corresponding codes are publicly available.
-Keywords: Newton-Kaczmarz method, piecewise-linear functions, training on disjoint datasets.
+The present paper introduces concurrency-driven enhancements to the training algorithm for the Kolmogorov-Arnold networks (KANs) that is based on the Newton-Kaczmarz (NK) method.
+As indicated by prior research, the NK-based training for KANs offers state-of-the-art performance in terms of accuracy and training time on relatively large datasets, significantly overtaking classical neural networks based on multilayer perceptrons (MLPs).
+Although some elements of the algorithm can be parallelised (in particular, evaluation of the basis functions’ values), a major limitation of the algorithm is the sequential application of the parameters’ updates, which has not been resolved up to now.
+However, substantial acceleration is achievable.
+Three complementary strategies are proposed in the present paper: (i) a pre-training procedure tailored to the NK updates’ structure, (ii) training on disjoint subsets of data, followed by models’ merging, not in the context of federated learning, but as a mechanism for accelerating the convergence, and (iii) a parallelisation technique suitable for execution on field-programmable gate arrays (FPGAs), which is implemented and tested directly on the device.
+All experimental results presented in this work are fully reproducible, with the complete source codes available online.
+Keywords: Kolmogorov-Arnold networks, Newton-Kaczmarz method, parallel training on disjoint datasets, FPGA.
 1  Introduction
-Since the publication of Liu et al.
-[1], which advocated spline-based univariate functions together with the ‘Adam’ or LBFGS optimisers for the training, these design choices have become dominant in the recent literature on Kolmogorov-Arnold networks (KANs).
-Several works have proposed alternative basis functions, aimed at reducing the computational overhead relative to neural networks, such as radial or alternative polynomial bases.
-However, these approaches remain conceptually similar — they rely on smooth basis functions [^1] Term ‘smooth’ is used loosely, implying the existence of some number of continuous derivatives; formally, the term of course means belonging to $C^{\infty}$ class.
-and typically require a relatively large number of calculations per each underlying function of KAN.
-Earlier alternatives, predating [1], include ReLU-type and piecewise-linear basis functions [2, 3], which were observed to be faster, but most subsequent research followed [1] as the principal reference point.
-It is important to emphasise that KANs were not invented in 2024 and originate from the pioneering work of Igelnik and Parikh [4] published in 2003.
-A historical overview of KANs’ development can be found in e.g.
+A Kolmogorov-Arnold network (KAN) is a non-linear regression model, i.e.
+a “black box” that converts an input vector to an output vector, alternative to a classical neural network — multilayer perceptron (MLP).
+There are important differences between KANs and MLPs, discussion of which the reader can find in abundant literature on this topic.
+Recent popularity of KANs can be attributed to [1].
+However, KANs were not invented in 2024 and originate from the pioneering work of Igelnik and Parikh [2] published in 2003.
+A brief historical overview of KANs’ development can be found in e.g.
 [3] and references therein.
-A number of recent studies have evaluated the federated learning performance of KANs relative to multilayer perceptrons (MLPs) [6, 7, 8, 9].
-These works also adopt the spline-based architectures and various optimisation methods.
-While improved accuracy for KANs is consistently reported, the computational cost is significantly higher.
-For example, the authors of [6] report training times of $314\,\text{s}$ for their KAN variant vs.
-$2.35\,\text{s}$ for an MLP on the same task.
-In the present paper, it is shown that such difference is not intrinsic to the KAN model class; rather, it arises from particular choices of basis functions and optimisation algorithms.
-For the piecewise-linear basis functions and the Newton-Kaczmarz optimisation scheme, this disparity is reversed, and highly-efficient KANs’ training is achieved while maintaining comparable accuracy.
-Moreover, in the present paper, for the first time, it is shown that the application of the federated-learning idea, i.e.
-training on disjoint subsets of a training dataset and merging the models, can significantly accelerate the training and represents a form of parallelisation of the training.
+To construct models based on KANs, one needs to make a number of choices.
+In particular, KANs contain basis functions (a rough analogy of them may be activation functions in MLPs), which can be chosen to be splines, Gaussians, piecewise-linear functions, other polynomials, etc.
+One then needs to choose a training method (also called ‘optimiser’), which fits the parameters of the model to the given dataset.
+Other important choices include the number of layers in the model (again, a rough analogy may be a number of layers in MLPs) and the number of basis functions within one layer.
+Following [1], the popular choices nowadays are the spline basis functions combined with ‘Adam’ or LBFGS training methods.
+However, the primary consequence of such choice is poor training time (along with other drawbacks discussed below).
+Earlier approaches [4, 5] advocated the use of piecewise-linear basis functions, which can achieve the same accuracy within a significantly shorter training time (for models that do not require the continuity of the derivatives of the outputs by the inputs).
+Furthermore, in [5] an efficient training algorithm based on the Kaczmarz method for solving linear systems of equations has been proposed; the algorithm has been improved in [3], where it has also been shown that the approach has a significant performance edge.
+The approach however relies on a series of sequential steps, parallelisation of which has not been resolved up to now.
+The aim of the present paper is to propose three strategies that improve the training time of an already efficient training algorithm for KANs based on the Newton-Kaczmarz method and proposed in [3].
+The strategies are (i) a new pre-training procedure, (ii) a concurrent training approach using disjoint subsets of the dataset together with the subsequent merging of the models, (iii) a parallelisation technique suitable for execution on field-programmable gate arrays (FPGAs).
+The paper aims to demonstrate the first implementation of KANs’ training on FPGAs.
+Strategies (i) and (ii) are general and are valid for any choice of the basis functions.
+The implementation on FPGAs however focuses on piecewise-linear basis functions.
 Although basis functions with continuous derivatives are essential for some applications, such as physics-informed modelling, they are not universally required.
 Moreover, the basis used during the training does not need to coincide with the basis used in the final model.
-In particular, a model may be trained efficiently using a piecewise-linear representation and subsequently upgraded to splines or other smooth bases with only minor post-training, thereby preserving accuracy while substantially reducing the training time.
+In particular, a model may be trained efficiently using a piecewise-linear representation and subsequently upgraded to splines or other bases with only minor post-training, thereby preserving accuracy while substantially reducing the training time.
 Thus, selecting piecewise-linear functions does not represent a move towards a more primitive model; instead, it enables faster optimisation while retaining full descriptive capabilities.
 The latter is further elaborated in [3].
-2  Approach
-A Kaczmarz-based training method for KANs with piecewise-linear basis functions was introduced by the authors in 2020 in a preprint later published as [3]; the approach was generalised to the Newton-Kaczmarz training method for KANs with arbitrary basis functions in 2023 in a preprint later published as [5].
-The reader is referred to the above publications for details; however, for completeness of the paper, a brief summary is provided below, assuming that the reader is familiar with KANs.
+The spline-based KAN implementation trained with ‘Adam’ and LBFGS methods from [1] has been rapidly superseded by FastKAN [6].
+While the improvements significantly reduce training time, the resulting performance remains well below the practical limits achievable with alternative approaches.
+Moreover, FastKAN users remain constrained by software-level design decisions, including dependency on specific runtime environments.
+These constraints raise questions regarding deployability and the integration of training procedures into non-Python-based applications.
+A notable property of KANs, independently observed by multiple authors [7, 8, 9, 10] and exploited in this work, is the possibility to merge models trained on disjoint datasets in a straightforward and efficient manner.
+In contrast, conventional neural networks generally lack an equally simple merging mechanism.
+To date, this property has been primarily investigated in the context of federated learning.
+The present work instead examines its use for concurrent training on disjoint data subsets.
+Existing work on FPGA implementations of KANs has primarily focused on hardware acceleration of inference rather than training.
+In particular, prior studies have taken models trained using floating-point arithmetic, converted them into FPGA-compatible representations using lookup tables (LUTs), block RAM (BRAM), and DSP resources, and evaluated the performance of the resulting forward-pass computations [11].
+The reported speedups, in some cases exceeding three orders of magnitude, reflect improvements in inference throughput relative to software-based forward-pass implementations.
+While these results demonstrate the feasibility and efficiency of FPGA-based inference for KANs, they do not address on-device training or end-to-end learning performance.
+The present work complements these efforts by extending FPGA utilisation beyond inference and towards accelerating the training process itself.
+2  The core concept of the approach
+A Kaczmarz-based training method for KANs with piecewise-linear basis functions was introduced by the authors in 2020 in a preprint later published as [5]; the approach was generalised to the Newton-Kaczmarz training method for KANs with arbitrary basis functions in 2023 in a preprint later published as [3], where the convergence was proven.
+The reader is referred to the above publications for details; however, for completeness of the paper, a brief summary is provided below.
 The basic building block of KAN is an operator that converts $n$-dimensional vector $\boldsymbol{y}$ into scalar $\hat{z}$.
-It can be called the discrete Urysohn operator or a generalised additive model:
+It can be called the discrete Urysohn operator [12] or a generalised additive model [13]:
 $\hat{z}=\sum_{j=1}^{n}g_{j}\left(y_{j}\right),$ (1)
-where $g_{j}$ are some functions. A layer of KAN consists of a set of such blocks, forming $m$-dimensional vector $\hat{\boldsymbol{z}}$, such that
+where $g_{j}$ are some functions.
+These vector-to-scalar mappings are referred to just as ‘blocks’ below.
+A layer of KAN consists of a set of such blocks, forming $m$-dimensional vector $\hat{\boldsymbol{z}}$, such that
 $\hat{z}_{i}=\sum_{j=1}^{n}g_{i,j}\left(y_{j}\right),\quad i=1,2,\ldots,m.$ (2)
@@
 where $\Delta\boldsymbol{z}=\boldsymbol{z}-\boldsymbol{\hat{z}}$ and $\Delta\boldsymbol{y}=\boldsymbol{y}-\boldsymbol{\hat{y}}$ are the residuals; matrix $J$ consists of partial derivatives, such that $J_{i,j}=\partial g_{i,j}/\partial y_{j}$, which are trivial to compute for piecewise-linear functions.
-3  Merging of the models
+3  Proposed improvements
+3.1  Pre-training
+The classical vector-input-scalar-output two-layer KAN is the following model:
+$\hat{z}=\sum_{j=1}^{n}A_{j},\qquad A_{j}=g_{j}\left(\sum_{l=1}^{p}h_{j,l}\left(x_{l}\right)\right),$ (8)
+where $A_{j}$ stands for one addend.
+The addends are initialised at random (for details see preceding papers by the authors).
+The novel pre-training concept consists in creating models consisting of groups of addends:
+$\hat{z}^{(1)}=\sum_{j=1}^{v}A_{j},\qquad\hat{z}^{(2)}=\sum_{j=v+1}^{2v}A_{j},\qquad\ldots\qquad\hat{z}^{(p/v)}=\sum_{j=p-v+1}^{p}A_{j}.\qquad$
+These models are trained concurrently, each approximating target $z$. After training, these models are assembled into original model (8) with the appropriate scaling (by $p/v$).
+In the case of a three-layer model, the main idea of the pre-training consists in training a classical (two-layered) model first, then disregarding the top layer and taking the intermediate variable values as the new inputs, using which another classical (two-layered) model is trained.
+This gives the initial approximation for three layers, which are then trained in a standard way.
+Each classical (two-layered) model can then be trained either from the initial random state or with the just described pre-training algorithm for the two layers.
+Analogously, pre-training for an arbitrary multi-layered version of the model can be done via a loop, where at each step the following is performed: two layers are trained, the outer layer is disregarded, the inner layer is used to create new input vectors.
+Each two-layer model can then be pre-trained as above.
+3.2  Merging of models
 The described approach iterates record-by-record through the training dataset. The hypothesis of the present paper is that such training can be parallelised using the algorithm as follows.
@@
 Thus, in the sequential case, each update is performed for the up-to-date model, while in the parallel case, threads starting from the second work with the ‘old’ model.
-4  Experimental part performed on laptop — strong scaling
-4.1  Baseline
+3.3  Parallelisation for FPGAs
+As seen from equation (2), each layer of KAN consists in evaluation of the independent functions with their subsequent summation.
+Since functions are independent, their evaluation during the forward pass can be performed concurrently, while successive layers should be processed sequentially due to data dependencies.
+This results in a small fixed number of sequential stages per layer, independent of the dimensionality of the involved vectors.
+Due to a small number of operations between synchronisation steps, such implementation will be inefficient on a regular multi-core x86-64 architecture.
+However, it is suitable for alternative hardware, such as FPGAs.
+When the training is performed, once the targets for each layer are calculated, updates of all functions of a given layer become independent.
+This means that they can be performed concurrently, exactly as for the forward pass.
+Consequently, the complete forward evaluation, calculation of the residuals for the layers, and parameters’ updates for one training record can be performed in a small fixed number of sequential stages, determined exclusively by the number of layers and not by the size of the layers themselves.
+Application-specific integrated circuits (ASICs) and FPGAs are well suited for such execution model.
+While practical implementations may deviate from the idealised concurrency assumptions and require more than the minimal number of cycles per record, the underlying structure remains highly amenable to hardware parallelism.
+Even on modest FPGA platforms with typical clock periods of the order of $10\,\mathrm{ns}$, the processing time of one record (forward pass and parameters’ updates) can be lower than that on modern x86-64 CPUs due to the described parallelisation mechanism.
+3.3.1  Implementation remarks
+Implementing computational routines on FPGAs and ASICs involves several well-known challenges.
+First, fixed-point arithmetic is preferred over floating-point arithmetic.
+The presence of a decimal point complicates hardware operations, whereas integer arithmetic is simpler and more efficient.
+Importantly, FPGA integers are not restricted to standard word sizes: if 32 bits are insufficient, arbitrary bit widths can be used without resorting to 64-bit representations.
+Second, division operations should be avoided whenever possible.
+Although division does not prohibit implementation, it may significantly degrade performance.
+Third, and most critically, computations expressed as clocked loops and nested if/else structures must be designed to execute a nearly constant number of operations regardless of the control-flow path.
+Since a KAN model is structured as a tree of simple blocks, all arguments and function values can be safely rescaled, including conversion to integer representations.
+By allowing arbitrary rescaling of the domain, $\Delta y$ from equation (3) is chosen as a power of $2$.
+As a result, division is replaced by a binary shift, while the offset becomes an integer obtained via a bitwise mask.
+The fractional part is then represented by an integer numerator, with the denominator applied later through a binary shift.
+Elements of matrix $J$ from equation (7) are used to compute the inner products without explicitly applying the denominator.
+The denominator is instead applied to the final result using a binary shift.
+Similarly, the learning rate is implemented as a binary shift; for example, using $1/64$ instead of $0.02$ does not lead to any noticeable degradation of the training.
+4  Numerical experiments
+The aim of this section is to demonstrate the proposed improvements of the training procedure using the following three examples:
+  • Det4: Prediction of determinants of random $4\times 4$ matrices.
+The inputs are $16$ elements of a matrix, each between 0 and $1$; the output is the determinant of the matrix.
+The training dataset contains $100\mathrm{K}$ records, the validation dataset contains $20\mathrm{K}$ records.
+  • Det5: Prediction of determinants of random $4\times 4$ matrices.
+The inputs are $25$ elements of a matrix, each between 0 and $1$; the output is the determinant of the matrix.
+The training dataset contains $10\mathrm{M}$ records, the validation dataset contains $2\mathrm{M}$ records.
+  • Tetra: Prediction of the areas of triangular faces of randomly-generated tetrahedra ($4$ faces).
+The inputs are $12$ coordinates of the vertices of the tetrahedra in 3D ($4$ vertices, $3$ coordinates for each), each between 0 and $1$; the outputs are $4$ areas.
+The training dataset contains $500\mathrm{K}$ records, the validation dataset contains $50\mathrm{K}$ records.
+Predictive performance is assessed using the Pearson correlation coefficient (in percent) between the real and the modelled output.
+Because the target distributions are strongly concentrated around their mean, small root mean square errors can arise even when a model produces near-mean random outputs.
+In this context, the correlation provides a more discriminative measure of meaningful predictive behaviour.
+4.1  Comparison test 1: Det4 example, laptop CPU/GPU
 The single-threaded implementation of the approach described in section 2 has already been extensively tested in the previous publication by the authors.
 It has been shown that it significantly outperforms state-of-the-art MATLAB’s neural networks.
-However, for completeness of the present paper, it is worth re-establishing a baseline result — showing a brief comparison between the considered single-threaded (sequential) implementation and MATLAB’s neural networks.
-Tests are conducted on a CPU (Intel(R) Core(TM) Ultra 7 155H, 16 cores) and on a GPU (Nvidia GeForce RTX 4050 Laptop).
+However, for completeness of the present paper, it is worth re-establishing a baseline result — showing a brief comparison between the considered single-threaded (sequential) implementation and MATLAB’s neural networks (version R2025b), FastKAN, and Keras (TensorFlow 2.20.0, Keras 3.12.0).
+The comparison is extended with the proposed multi-threaded improvements: use of disjoints and use of pre-training for KANs.
+The tests were conducted on a CPU (Intel(R) Core(TM) Ultra 7 155H, 16 cores) and on a GPU (Nvidia GeForce RTX 4050 Laptop).
 Obviously, one cannot objectively compare a CPU code with a GPU code, as they run on completely different hardware.
@@
 whether performance gain of MATLAB’s neural networks when switching to a GPU is sufficient to overtake the authors KANs’ implementation on such particular hardware.
-For the tests, synthetic data is generated.
-The inputs are elements of $4\times 4$ matrices and the outputs are the determinants of the matrices.
-The training dataset contains $100\mathrm{K}$ records; the validation dataset contains $20\mathrm{K}$ records.
-Random data is generated each execution of the code.
-The accuracy metric is the Pearson correlation coefficient (in percent) between the real and the modelled output.
-This is just a demonstrative example.
-Therefore, only one configuration of MATLAB’s neural network is used, with $5$ hidden layers, each consisting of $10$ neurons (i.e.
-the total of $621$ parameters per model); only one KAN configuration in the authors’ code is used, with $70$ addends (classical two-layer architecture), $3$ points per inner function, and $25$ points per outer function (i.e.
-the total of $5,110$ parameters per model).
-Readers who are interested in a more detailed comparison with varying MLP and KAN architectures are referred to the previous paper [5].
-The results are shown in table 1, where it is seen that MLPs are outperformed by the authors’ implementation of KANs.
-When switching to a GPU, MATLAB’s neural networks are approximately $3$ times faster (on the particular tested hardware), but still have slightly lower accuracy and longer training time than KANs.
-MLP, CPU MLP, GPU KAN
-Pearson (%) $94.6\pm 1.4$ $95.0\pm 0.9$ $97.1\pm 0.1$
-time (s) $32.7\pm 0.9$ $8.6\pm 0.3$ $5.9\pm 0.6$
-Table 1: Baseline accuracy and performance tests for modelling of the determinants of random $4\times 4$ matrices conducted for MATLAB’s neural networks and the authors’ KANs’ implementation (single-threaded, C++ implementation for CPU).
-The values are expectations and standard deviations for 88 executions.
-4.2  Scaling test
+The results for the Det4 example are summarised in table 1.
+The reported values are the means and the standard deviations, obtained using $8$ independent executions, where the datasets and the models are initialised at random.
+The accuracy is of course measured on the validation dataset.
+The results are consistent across MATLAB, FastKAN, and Keras, including their GPU-accelerated versions, and closely match the sequential C++ KANs’ implementation run on CPU.
+The extended version, which incorporates pre-training and training on disjoint subsets, achieves a speedup of approximately $30$ times over the baseline CPU tests and about $7$ times over the corresponding GPU tests.
+Table 1: The accuracy and performance tests for Det4 example. Comparison of MATLAB’s neural networks, FastKAN, Keras, and the authors’ C++ KANs’ implementation.
+Pearson (%) time (s)
+MATLAB CPU $94.6\pm 1.4$ $32.7\pm 0.9$
+MATLAB GPU $95.0\pm 0.9$ $8.6\pm 0.3$
+FastKAN CPU, first config. $97.3\pm 1.2$ $65.7\pm 4.7$
+FastKAN GPU, first config. $97.5\pm 1.0$ $19.6\pm 2.1$
+FastKAN CPU, second config. $95.6\pm 1.1$ $44.3\pm 3.6$
+FastKAN GPU, second config. $95.2\pm 0.9$ $9.2\pm 1.7$
+Keras CPU, first config. $97.0\pm 0.0$ $30.6\pm 2.1$
+Keras GPU, first config. $97.0\pm 0.0$ $7.5\pm 1.4$
+Keras CPU, second config. $94.0\pm 0.0$ $31.2\pm 1.9$
+Keras GPU, second config. $94.0\pm 0.0$ $6.8\pm 0.7$
+C++ sequential $97.2\pm 1.1$ $6.57\pm 0.90$
+C++ disjoints $97.4\pm 0.9$ $2.80\pm 0.12$
+C++ disjoints and pre-training $97.5\pm 2.1$ $0.98\pm 0.09$
+C++ disjoints and pre-training run on Linux $97.4\pm 1.4$ $0.70\pm 0.02$
+Of course, CPU and GPU implementations run on completely different hardware architectures, and their execution times are therefore not directly comparable.
+Nevertheless, this comparison is included because the experiments were conducted on standard laptop hardware, and switching between CPU and GPU execution requires only a minimal change in the execution script.
+In practice, many users take advantage of this convenience and the resulting performance gains.
+Moreover, the authors have been repeatedly asked about such comparisons in technical discussions, which further motivated their inclusion in this paper.
+MATLAB’s neural network configuration.
+Only one configuration of MATLAB’s neural network is used, with $5$ hidden layers, each consisting of $10$ neurons (the total of $621$ parameters per model), and the activation functions being hyperbolic tangents.
+Readers interested in a more detailed comparison with varying MLP architectures are referred to the previous paper [3].
+The difference between the previous publication and results demonstrated here is that the single-precision floating point format is used in MATLAB in the tests below (compared to the double-precision format used previously), which significantly accelerates its performance.
+FastKAN requires installation of a dedicated software framework and its associated runtime environment.
+The package includes a single example provided by the developers — MNIST dataset, consisting of $60\mathrm{K}$ training records and $10\mathrm{K}$ validation records, with $784$ input features and $10$ output targets.
+The default model used in the unit test has two layers containing $64$ and $10$ blocks.
+Therefore, the number of the underlying functions in the inner and the outer layers are $64\times 784$ (one underlying function per one input per one block) and $64\times 10$ (same as for the previous layer, but the number of the inputs for the second layer is of course $64$), respectively.
+When the number of splines (basis functions) per function is unspecified, the default value of $64$ is used.
+Each spline (basis function) enters the model with one coefficient (parameter); therefore, the default model has the total of $64\times(64\times 784+64\times 10)=3{,}252{,}224$ parameters.
+This drastically contrasts with the number of records in the dataset — only 60,00060{,}000, which is exactly the number of equations, from which the parameters are determined.
+Even first-year undergraduate students are taught that models should be constructed such that the number of parameters should be at least smaller than the number of equations.
+The authors do not understand the reason for such badly designed default example, which also significantly misleads the users in terms of the sizes of the networks to be used.
+One possible explanation is that since FastKAN uses parallel processing, increasing the model size can paradoxically reduce training time in some cases, as more parameters allow the optimiser to reach local minimum faster.
+FastKAN configuration.
+Two configurations of FastKAN’s network are used.
+The first configuration has four layers with $512$, 88, $16$, and $1$ blocks; each underlying function is represented by $10$ splines; the total of $10\times(16\times 512+512\times 8+8\times 16+16\times 1)=124{,}320$ parameters.
+The second configuration has four layers with $256$, 88, $16$, and $1$ blocks; each underlying function is represented by $8$ splines; the total of $8\times(16\times 256+256\times 8+8\times 16+16\times 1)=50{,}304$ parameters.
+The first configuration has the larger number of parameters than training records — this of course should not be done in a general case (and the authors understand that well).
+Keras’s neural network configuration.
+Two configurations of Keras’s neural network is used, with ReLU activation functions.
+The first configuration has $3$ hidden layers, consisting of $256$, $32$, and $24$ neurons (the total of 13,39313{,}393 parameters per model).
+The second configuration has $3$ hidden layers, consisting of $64$, $32$, and $32$ neurons (the total of 4,2574{,}257 parameters per model).
+Authors’ KAN configuration.
+Only one KAN configuration in the authors’ code is used, with $2$ layers (classical architecture), consisting of $70$ and $1$ blocks, $3$ points per function of the first block, and $30$ points per function of the second block (the total of 5,4605{,}460 parameters per model).
+Four scenarios are considered: (a) purely sequential execution without any concurrency, (b) training on three disjoint subsets with simple merging via averaging, (c) the same as previous with the addition of pre-training, (d) the same as previous run on a Linux operating system instead of Windows.
+4.2  Comparison test 2: Tetra example, laptop CPU/GPU
+The results for the Tetra example are summarised in table 2.
+It is seen that when trained to the same accuracy, parallel implementation of KANs is about $2.8$ times faster than the sequential and two orders of magnitude faster than MLP training on CPU.
+MATLAB’s neural networks with $4$ hidden layers and $80$ neurons per layer ($20,804$ model parameters) are trained.
+Three-layered KANs are trained, the first layer consisting of $70$ operators with $3$ points per function, the second layer consisting of $30$ operators with $18$ points per function, the third layer consisting of $4$ operators with $22$ points per function (the total of $42,960$ model parameters).
+Table 2: The accuracy and performance tests for Tetra example. Comparison of MATLAB’s neural networks and the authors’ C++ KANs’ implementation.
+Pearson (%) time (s)
+MATLAB CPU $98.6\pm 0.7$ $98.2\pm 0.9$ $98.9\pm 0.3$ $97.9\pm 0.4$ $1126.4\pm 69.5$
+MATLAB GPU $98.7\pm 0.5$ $98.3\pm 0.7$ $98.2\pm 0.8$ $98.3\pm 0.9$ $76.9\pm 1.2$
+C++ KAN sequential $8.6\pm 0.6 99$$9.0\pm 0.4$ $98.1\pm 0.2$ $98.5\pm 0.3$ $29.2\pm 3.2$
+C++ KAN disjoints, 6 threads $98.5\pm 0.9$ $98.1\pm 0.7$ $98.6\pm 0.3$ $98.3\pm 0.4$ $10.5\pm 1.3$
+4.3  Scalability test 1: strong scaling, laptop CPU
 In this section, the main hypothesis of the paper that parallelisation can be implemented by training on disjoint batches across multiple threads is benchmarked.
@@
 As usual, the speedup is measured as $S=t(1)/t(n)$, where $t(1)$ is the time required to complete the work on $1$ thread and $t(n)$ is the time required to complete the same work on $n$ threads.
-In the single-threaded setting for predicting determinants of $4\times 4$ matrices, approximately $10$ passes through the training dataset (epochs) are required to reach the Pearson correlation of at least $97\%$.
+In the single-threaded setting for Det4 example, approximately $10$ passes through the training dataset (epochs) are required to reach the Pearson correlation of at least $97\%$.
 This corresponds to processing $10^{6}$ records, and this value is used as the fixed workload for all parallel configurations.
@@
 Such loss in accuracy can be compensated of course by increasing the number of rounds.
-For example, using $6$ threads, $20$ rounds, and the batch size of $16,666$ (twice larger computational work than in the table) yields the Pearson correlation of $97.0\%$, while completing in $2.22$ seconds — $2.4$ times faster than for $1$ thread.
+For example, using $6$ threads, $20$ rounds, and the batch size of 16,66616{,}666 (twice larger computational work than in the table) yields the Pearson correlation of $97.0\%$, while completing in $2.22$ seconds — $2.4$ times faster than for $1$ thread.
+Table 3: The strong scaling test for the authors’ C++ KANs’ implementation.
 threads rounds batch size Pearson (%) time (s)
-1 10 100,000 96.8 5.4
+1 10 100,000 96.8 5.41
 2 10 50,000 96.7 2.58
@@
 6 40 4,166 94.6 1.19
-Table 2: Accuracies and execution times of the code on a laptop for the example of matrix determinants — strong scaling test.
-4.3  Second example
-The second example is not strictly a scaling test, as the computational workload varies, but it serves as a benchmark of parallel training to a fixed accuracy.
-The training is done on synthetic data, where the outputs are the areas of triangular faces of randomly-generated tetrahedra ($4$ faces), and the inputs are the coordinates of the vertices of the tetrahedra in 3D ($4$ vertices, $3$ coordinates for each).
-Number of the training records is $500\mathrm{K}$, number of the validation records is $50\mathrm{K}$.
-Three-layered KANs are trained, the first layer consisting of $70$ operators with $3$ points per function, the second layer consisting of $30$ operators with $18$ points per function, the third layer consisting of $4$ operators with $22$ points per function (i.e.
-the total of $42,960$ model parameters).
-MATLAB’s neural networks are used for the comparison with $4$ hidden layers and $80$ neurons per layer ($20,804$ model parameters).
-The results are shown in table 3, where it is seen that when trained to the same accuracy, parallel implementation of KANs is about $2.8$ times faster than the sequential and two orders of magnitude faster than MLP training on CPU.
-Pearson (%) for outputs time (s)
-MLP, CPU $98.6\pm 0.7$ $98.2\pm 0.9$ $98.9\pm 0.3$ $97.9\pm 0.4$ $1126.4\pm 69.5$
-MLP, GPU $98.7\pm 0.5$ $98.3\pm 0.7$ $98.2\pm 0.8$ $98.3\pm 0.9$ $76.9\pm 1.2$
-KAN sequential $8.6\pm 0.6 99$$9.0\pm 0.4$ $98.1\pm 0.2$ $98.5\pm 0.3$ $29.2\pm 3.2$
-KAN parallel, 6 threads $98.5\pm 0.9$ $98.1\pm 0.7$ $98.6\pm 0.3$ $98.3\pm 0.4$ $10.5\pm 1.3$
-Table 3: Baseline accuracy and performance tests for modelling of the areas of faces of random tetrahedrons conducted for MATLAB’s neural networks and the authors’ KANs’ implementation.
-5  Experimental part performed on cluster — weak scaling
+4.4  Scaling test 2: weak scaling, HPC cluster
 To assess the scalability of the code, the tests were run on an HPC cluster.
 All jobs were run on a single node equipped with two Intel(R) Xeon(R) Platinum 8358 CPUs ($32$ cores per CPU) and $256\mathrm{GB}$ RAM.
-The tests consisted in training KANs on datasets with the inputs being elements of $4\times 4$ or $5\times 5$ matrices and the outputs being the determinants of the matrices.
+Two examples are considered: Det4 and Det5.
 In this section, the weak scaling is tested: the number of threads was kept proportional to the total number of passes through the data (epochs), i.e.
@@
 The results are shown in table 4, where it is seen that the scaling is very close to linear.
-For the example of $4\times 4$ matrix determinants, the efficiency is above $93\%$ when up to $32$ threads are used.
-For the example of $5\times 5$ matrix determinants, the efficiency is above $95\%$ when up to $16$ threads are used and only drops down to $83\%$ when $32$ threads are used (and even further for $64$ threads).
-Such noticeable decline in the latter case is due to a large number of model synchronisations and large model sizes for $5\times 5$ example.
-In the example of $4\times 4$ matrix determinants, the datasets consisted of $100\mathrm{K}$ training records, the models consisted of $70$ addends, with $3$ points per inner function and $25$ points per outer function (i.e.
-the total of $5,110$ parameters per model).
+The values are averages of three programme runs.
+For Det4 example, the efficiency is above $93\%$ when up to $32$ threads are used.
+For Det5 example, the efficiency is above $95\%$ when up to $16$ threads are used and only drops down to $83\%$ when $32$ threads are used (and even further for $64$ threads).
+Such noticeable decline in the latter case is due to a large number of model synchronisations and large model sizes for Det5 example.
+In Det4 example, two-layer KANs are constructed, with $70$ blocks in the first layer and $3$ points per function, $1$ block in the second layer and $25$ points per function (the total of 5,1105{,}110 parameters per model).
 The batch size was taken to be $100\mathrm{K}$ records, i.e.
@@
 $10$ synchronisations (of $n$ models, where $n$ is the number of threads) are made within each execution of the code.
-In the example of $5\times 5$ matrix determinants, the datasets consisted of $10\mathrm{M}$ training records, the models consisted of $160$ addends, with $3$ points per inner functions and $30$ points per outer function (i.e.
-the total of $16,800$ parameters per model).
+In Det5 example, two-layer KANs are constructed, with $160$ blocks in the first layer and $3$ points per functions, $1$ block in the second layer and $30$ points per function (the total of 16,80016{,}800 parameters per model).
 The batch size was taken to be $50\mathrm{K}$ records.
 The jobs consisted of $200$ batches.
+Table 4: The weak scaling test for the authors’ C++ KANs’ implementation.
 threads 11 22 44 88 $16$ $32$ $64$
-time (s), $4\times 4$ $6.64$ $6.80$ $6.82$ $6.82$ $7.11$ $7.04$ $7.79$
-time (s), $5\times 5$ $217.02$ $220.24$ $226.07$ $222.82$ $227.68$ $260.99$ $282.96$
-Table 4: Execution times of the code on HPC cluster for examples of matrix determinants — weak scaling test. The values are averages of three programme runs.
-6  Concurrent-serial alternations of Newton-Kaczmarz
-The conceptual description of the Newton-Kaczmarz algorithm provided in section 2 highlights its alternating concurrent-sequential structure.
-Each layer maps an input vector to an output vector through multiple independent operations that may be executed concurrently.
-However, the outputs of all concurrent computations must be synchronised before proceeding to the next layer.
-The backward updates follow the same pattern — a chain of potentially concurrent steps separated by unavoidable sequential dependencies.
-These serial steps between independent computations represent a significant obstacle for concurrent training.
-Neither CPU nor GPU kernels are well suited for such patterns.
-This limitation can be illustrated using the classical vector-input-scalar-output two-layer KAN:
-$\hat{z}=\sum_{j=1}^{n}A_{j},\qquad A_{j}=g_{j}\left(\sum_{l=1}^{p}f_{j,l}\left(x_{l}\right)\right),$ (8)
-where $A_{j}$ stands for one addend, and each addend is itself a small tree mapping inputs $x_{l}$ to a scalar.
-All $n$ addends must be computed in order to obtain the residual $\Delta z$, and these computations can be executed concurrently.
-This residual, divided evenly among the addends and scaled by the learning rate, is the only quantity required for updating all functions $g_{j}$ and $f_{j,l}$.
-In the piecewise linear model, evaluating a function requires only two operations (one division and one multiplication), which is far too little computation for concurrency on CPUs or GPUs to provide any advantage.
-The synchronisation required after all $n$ addends are computed often takes longer than the time saved through parallelism.
-However, this pattern is well suited for customised hardware, such as field-programmable gate arrays (FPGAs) or application-specific integrated circuits (ASICs).
-A CPU-based simulation demonstrating a custom synchronisation barrier designed by the authors is available online [^2] https://github.com/andrewpolar/detfast, but it serves only as a conceptual demonstration; synchronising at each training record introduces substantial overhead.
-The example implemented in the code is training a model for determinates of $5\times 5$ matrices.
-Training MLPs for the latter requires several hours [5], whereas the CPU-based KAN code completes it in a few minutes.
-An FPGA or ASIC implementation could reduce this further to seconds, because each addend can be mapped to a dedicated hardware processing block.
-In such customised hardware, concurrency is not emulated through threads or kernels but is realised through physical parallelism, where synchronization and data flow are governed directly by wiring rather than by software-controlled barriers.
-6.1  Pretraining
-Structure (8) motivates a pretraining strategy, in which groups of addends are trained concurrently over the entire dataset to predict an equal share of target $z$.
-For the example of determinants of $5\times 5$ matrices and a dataset of $10\mathrm{M}$ records, such pretraining produces a model with the Pearson correlation near $70\%$, providing a strong initialisation for the main training phase.
-When pretraining is combined with the concurrent-serial alternations, the full training process for $5\times 5$ determinants completes in just over $2$ minutes on the CPU used for the baseline examples.
+time (s), Det4 $6.64$ $6.80$ $6.82$ $6.82$ $7.11$ $7.04$ $7.79$
+time (s), Det5 $217.02$ $220.24$ $226.07$ $222.82$ $227.68$ $260.99$ $282.96$
+4.5  Testing of the implementation on FPGA board
+The implementation was run on Digilent Nexys A7-100T (Xilinx Artix-7) board.
+Two simple datasets are used: prediction of the Euclidean norms for random vectors of dimensions $4$ and $16$.
+The training dataset consists of $1024$ records, processed in $3$ cycles per record over $2$ epochs, for a total of $6144$ cycles.
+At a base clock frequency of $100\,\mathrm{MHz}$, the total execution time is $0.06144\,\mathrm{ms}$.
+The same number of latches is used for both the $4$-function and $16$-function models, as all forward-pass and update computations are performed concurrently.
 5  Conclusions
-The results demonstrate that KANs’ training can indeed be accelerated by iteratively training copies of the model on disjoint datasets (batches) and merging them.
-The speedup ratios between sequential and concurrent KAN training typically fall in the range of 3344, meaning that acceleration is possible, but there are some limits.
-At the same time, even the sequential KAN implementation by the authors consistently outperforms MLPs on CPUs and, remarkably, also on a GPU.
-While direct CPU-GPU comparisons are formally inappropriate due to differences in hardware, these results are included for reference, taking into account that the GPU used for the benchmarking can be considered to be a typical laptop GPU (it has been noticed by the authors that some researchers have unrealistic expectations from GPU implementations).
-Since methodology adopted in the authors’ code is FPGA-friendly, the authors believe that this direction is a perspective research capable to reduce hours into seconds of processing time for very large models and datasets.
-MATLAB scripts [^3] http://openkan.org/downloadskan/matlabscripts.zip, sequential implementation [^4] https://github.com/andrewpolar/kankan-8, and parallel implementation [^5] https://github.com/andrewpolar/meanmodel are available online.
+The test results demonstrate strong performance consistency across MATLAB, FastKAN, and Keras, while the authors’ C++ KANs’ implementation delivers significantly faster training times.
+Even purely sequential implementation remains competitive with GPU-accelerated versions of the other software on a ‘standard’ laptop with CUDA support, and concurrent implementation achieves up to a $7$ times speedup.
+Importantly, these benchmarks remain far from the theoretical performance limits.
+The computational flow of the considered KANs’ training method is exceptionally well-suited for hardware-supported concurrency available in FPGAs and ASICs, providing substantial potential for further acceleration.
+In such a hardware context, performance boundaries are determined not by software efficiency but by the size and capabilities of the hardware itself.
+Given ongoing industrial advances, KAN models trained using the proposed technique represent a highly promising avenue for scalable and high-performance AI.
+Beyond raw training speed, maintainability and deployability are crucial considerations.
+The demonstrated C++ implementations are concise (300300700700 lines), avoid third-party libraries, and do not rely on language-specific features, making them straightforward to port to other programming environments and integrate into larger applications.
+This combination of performance, portability, and scalability positions such KANs’ training as a practical and impactful approach for future AI development.
+MATLAB scripts [^1] http://openkan.org/downloadskan/matlabscripts.zip, a sequential C++ implementation [^2] https://github.com/andrewpolar/kankan-8, a parallel implementation [^3] https://github.com/andrewpolar/meanmodel, and register-transfer level (RTL) code [^4] https://github.com/andrewpolar/fpga for FPGAs are available online.
+The readers can also find useful discussions and simplified explanations of the presented concepts on the authors’ website [^5] http://openkan.org.
+Acknowledgements
+MP acknowledges funding from Research England’s “Expanding Excellence in England” (E3) fund via the “Multi-scale Multi-disciplinary Modelling for Impact” programme (M34Impact).